OpenStudy (anonymous):

18+22+26+30+...+50=? Find the sum of the given arithmetic series.

OpenStudy (anonymous):

hah ill let someone else help you with that;)

OpenStudy (swag):

@aminah.loves17 Please do not post answers like that if you will not help them.

OpenStudy (jwheele1):

is this a pattern question? Or do we know what the dots are supposed to be?

OpenStudy (anonymous):

its nothing bad so chill

OpenStudy (jwheele1):

ok well...d = 4

OpenStudy (jwheele1):

a = 18

OpenStudy (jwheele1):

Skip to one minute and 48 seconds in http://www.youtube.com/watch?v=UHkueFmPC6s&feature=player_embedded#!

OpenStudy (jwheele1):

ill do it the long ugly way: 18+22+26+30+34 + 38 + 42+46+50

OpenStudy (jwheele1):

18+22+26+30+34 + 38 + 42+46+50 40 + 56 + 72 + 88 + 50

OpenStudy (jwheele1):

40 + 56 + 72 + 88 + 50 96 + 160 + 50

OpenStudy (ajprincess):

use the formula \[a_n=a+(n-1)d\] to find n. Here a=18, d=4 50=18+(n-1)4 50=18+4n-4 50=14+4n 50-14=4n 36=4n 36/4=n 9=n Nw use the formula \[s=\frac{n(a+l)}{2}\] where s-sum, n=9, a=18 and l=50

OpenStudy (jwheele1):

96 + 160 + 50 256 + 50

OpenStudy (jwheele1):

306

OpenStudy (jwheele1):

@ajprincess what answer do you come up with?

OpenStudy (ajprincess):

the same 306. @jwheel1

OpenStudy (jwheele1):

cool

OpenStudy (jwheele1):

n = 9 means that there are 9 total numbers in the sequence fyi

OpenStudy (jwheele1):

so your formula is that the sum of the sequence (s) =/is the number of items (9) times (the difference pattern (4) + one additional step to the right in the sequence) and all of that is divided by 2 (and im not sure why, lol)

OpenStudy (anonymous):

\[Sn = \frac{ n ( a+ l) }{ 2 }\] \[an = a + (n-1)d\]